# Planck stars, White Holes, Remnants and Planck-mass quasi-particles The quantum gravity phase in black holes’ evolution and its manifestations

### v3

@@ I Introduction
Three aspects of this scenario are particularly appealing. (a) It offers a natural solution to the black hole information ‘paradox’. (b) It is in principle, and perhaps even in practice, directly testable Perez2023 . And (c) it provides a natural candidate for dark matter that does not require any new physics (such as new fields, particles, or modifications of the field equations): just general relativity and its quantum properties Rovelli2018f . This scenario is compatible and possibly corroborating the general idea that primordial black holes could play a key role for explaining dark matter, see carr_primordial_2020 ; green_primordial_2021 .
@@
In the first, we discuss the bounce of a collapsing matter distribution (Section II.1), namely the ‘Planck Star’ phenomenon, and the external metric of the global black hole dynamics (Section II.2). A side section discusses the counter intuitive classical geometry of white holes, because this play a role in global picture (Section II.6). This part ends presenting the current state of the calculations of the black-to-white transition probability Christodoulou2018 .
@@ I.1 A sketch of the scenario
Perhaps the most interesting discovery of the last decades is the abundance and variety of the astrophysical object we now call black holes. Current direct and indirect observations of these objects are all well accounted for by classical GR. But classical GR is insufficient to account for the phenomena that happen in their high-curvature regions, where genuine quantum gravity effects are most likely non-negligible. Therefore GR does not provide a reliable global picture of the dynamics of these objects.
@@
The internal geometry of these objects is a remarkably precise realization, fully consistent with classical GR, of the ‘cornucopia’ intuition considered in the 1980’s and later abandoned on the basis of arguments that (as we shall discuss below) are not anymore cogent Banks1992b ; Banks1993 .
There is a well known instability affecting eternal and macroscopic white holes: they can easily fall into a black hole as in the process described by the classical Kruskal geometry. On the other hand, there are arguments indicating that a Plank-mass white hole can be stabilized by quantum gravity Rovelli2018f , by shifting it into a quantum superposition of black and white geometries. As we show below, the key of this stabilization is the LQG area gap, because of which there is likely no black (or white) hole with a mass smaller than the Planck mass.
Hence the scenario predicts the existence of quasi-stable remnants of a known mass, determined by the LQG area gap. A large number of these, of primordial or possibly even pre-big-bang origin, would behave precisely as dark matter. The possibility of direct detection of these object has been explored in Perez2023 and we review it below.
Once quantum gravity phenomena are taken into account, the notion itself of horizon is necessarily different than the one relevant for the classical theory. Event horizons looses their relevance, because they are likely not to form at all. The Hawking radiation drags the dynamical evolution of the horizon into regions where the Einstein equations do not hold anymore. Accordingly, the discussion on the so called black-hole information paradox needs to be reconsidered.
@@
While the scenario illustrated takes fully into account quantum gravitational phenomena in the high curvature region, it assumes that classical GR and quantum field theory on curved spacetimes are good approximations outside these regions. Contrary to what sometimes claimed, GR and quantum field theory taken together, yield no inconsistencies in the entire region where quantum gravity can be neglected. As argued in detail below, the common claim that inconsistencies appear already at the Page time depends on strong versions of the holographic hypothesis Rovelli2017a , or from confusion between ADM mass and Bondi mass, between thermodynamic and von-Neumann entropies, or between event and dynamical horizons. On the contrary, the scenario considered here is only based on conventional GR and conventional quantum ideas.
@@ I.2 The domain of validity of classical gravity
The region of the onset of quantum gravitational phenomena can thus be organized into three sub regions Dambrosio2021 (see Figure 1 ):
@@
which is colossal for a macroscopic black hole. Notice that the locus of the onset of quantum gravity is not ‘at a point’, but rather ‘at a time’. It is not ‘somewhere’ but rather, ‘at some time’. Which in relativistic terms means in a set of space-like related regions.
The fact that the various onsets of quantum gravity are spacially related implies by locality that what happens in a region of this locus is causally disconnected, hence independent from what happens in other regions. To imagine that quantum gravity could violate this macroscopic causal disconnection is in gross contradiction with what we know about reality, not supported or suggest by anything we know, as far as we can see. To understand the physics of the end of a black hole, we have to understand the quantum evolution of each of these three regions independently.
@@ II.1 Planck stars
Let us start by studying what happens in the interior of a collapsing star, after it enters its horizon, when it reaches Planckian density. The simplest modeling of a collapsing matter distribution is provided by the Oppenheimer-Snyder model Oppenheimer:1939ue : a spherically symmetric pressure-less homogeneous matter distribution free-falling under its own weight. Assuming the matter distribution distribution to start at rest in the distant past, the metric inside a such a matter distribution distribution can be written in co-moving and proper time coordinates $(T,R)$ as
@@
where the overdot means differentiation with respect to $T$, which coincides with the proper time. Eq. (6) can be solved:
@@
where $A=3/(2\pi\rho_{c})$ is a parameter of Planckian value. In units where $G=c=1$, the constant is $A\sim\hbar\sim m_{Pl}^{2}$. The last equation can be integrated, giving
$a(T)=\left(\frac{9mT^{2}+Am}{2}\right)^{1/3}\,.$ (10)
Notice that $a(T)$ is positive for the whole range $T\in[-\infty,\infty]$: it decreases for $T<0$, reaches a minimum $a_{0}=\sqrt[3]{Am/2}$ for $T=0$ and then increases for $T>0$. See Fig. 2. This is the characteristic bounce of loop quantum cosmology. This result shows that the line element in eq. (5) is well defined everywhere, without singular collapse: the stars reaches a maximal density, where it is called a ‘Planck star’, and then bounces and expands.
@@
To get a feeling of the physics of the bounce, we can rewrite (8) in the form
@@ II.2 Black-to-white transition
The maximal extension of this geometry has (i) two asymptotic regions and (ii) a time-like singularity in a high curvature region. These two features make it an unlikely candidate for describing the actual physics of realistic black holes. But a surprising result on spherically symmetric solutions of the Einstein’s field equations, solve both these difficulties. This is illustrated in the next section.
@@ II.3 The exterior metric
The Birkhoff’s theorem is sometimes presented as stating that the only spherically symmetric spacetime compatible with the vacuum Einstein’s field equations is the Kruskal geometry. At a sufficient distance from a black hole, we expect quantum effects to be negligible and therefore the theorem to hold. Hence we might expected that whatever happens where quantum effects are negligible must form a subset of a Kruskal geometry. In the Kruskal metric, there is an anti-trapped region, but it is before, not after the trapped region. Hence there seem no way to join the metric described above to a single external asymptotic region. But the above is wrong. The reason is that Birkhoff’s theorem is local, not global: it states that a spherically symmetric solution of the vacuum Einstein field equations is locally, and not necessarily globally, isometric to the Kruskal metric. This seems a red hair, but has momentous consequences for understanding the dynamics of quantum black holes.
@@
Figure 4: The spacetime discovered in Haggard2014 , describing the collapse of a null shell into a black hole and its bounce out a white hole (left); and the way it can be locally isometrically mapped onto the Kruskal spacetime (right). The pink region is an exact solution of the classical Einstein equations.
@@
This same idea solves the two difficulties mentioned at the end of last section, providing a spacetime with a single asymptotic region and no singularities for the bouncing star. This can be constructed by cutting and gluing the maximal extension of the metric defined by eqs. (13) and (16), as follows.
@@
Next, delete the entire spacetime region enclosed within the blue line, and glue its two space-like portions as in Figure 6. This is clearly possible since they are both constant Schwarzschild time surfaces.
@@
The resulting spacetime has a single asymptotic region, is everywhere locally isomorphic to the maximal extension of the metric (13-16), and has a hole, depicted as the $B$ diamond of Figure 6. In Han:2023wxg it is shown that the $B$ region admits a regular Lorentzian metric compatible with rest of the spacetime geometry. The $B$ region and its physics will be discussed later on. For the moment, let us focus on the rest of the spacetime.
@@
The first term on the right hand side can be interpreted as the back and forth travel time for the light from the observer to the hole. The second term is the standard relativistic correction. The last term is independent from the radius and represents a property of the transition itself. The quantity $\delta$ can be measured by distant observers by measuring $T$ at different values of $R$. A little geometry shows that it depends on the choice of $\beta_{L}$ as follows: $r=2m+\delta$ is the radius at which the constant Schwarzschild-time surface passing by $\beta_{L}$ intersects the surface of the falling star. This is clearly very close to the collapse (the point where the surface enters entirely in its horizon), if the lifetime of the hole is long. Thus
@@
can be taken as a definition if the duration of the quantum process, as observed from infinity. In our construction, it is determined by how the geometry has been constructed cutting and gluing. Operationally, it is the observed duration of the process from very large distance. Physically, it is determined by the quantum theory of the geometry of the $B$ region, to which we now turn.
@@ II.4 The Boundary Region
To address this question we need to step back and think more precisely what we are doing. First, let us remind once more that for the moment we are disregarding the dissipative effects that break time reversal invariance. These will be studied in the next section. Then, consider the fact that we are describing a quantum process. There are different ways of describing quantum processes. In some regimes, these can be described as corrections to the classical equations of motion. This is not viable for the quantum transition of the horizon, which is a non perturbative process. One cannot describe quantum tunnelling as a simple $\hbar$ corrections to classical trajectories. Tunnelling is non analytical in $\hbar$.
Another possibility is to describe the evolution of a quantum state. Quantum states that approximate classical configurations, such as coherent states, spread. We could therefore consider the full quantum state of the geometry after the actual quantum process as a quantum states, namely a quantum superposition of geometries. This is analogous to describing nuclear radioactivity in terms of a wave function of the emitted particle widely spread in space and time. Correct, but incomplete. What we observe in nuclear radioactivity is not the wave function of the emitted particle, widely spread in space and time. Rather, it is a specific space and time location where the particle is detected, say by a Geiger counter.
@@
That is, we can compute the transition amplitude for the geometry to make a quantum transition from its value in the past boundary of $B$ to its future boundary (as opposite to not yet making it, as in radioactivity). Notice that the geometry (both intrinsic and extrinsic) of these two boundaries is known from the construction of the above section. We have simply to compute the LQG transition amplitudes between quantum states approximating these geometries.
The covariant formulation of LQG is precisely formulated to do this calculation, using approximations provided by truncations of the degrees of freedom. The calculation has been performed, under some drastic simplifications by Christodoulou and D’Ambrosio using analytical techniques in Christodoulou2018 ; Dambrosio2021 ; Christodoulou2023 . The result has been confirmed numerically in Frisoni2023 and Han2024 . Numerical calculations are currently developed to go beyond these approximations. The way these calculations are performed is summarized in the next section.
@@ II.5 The LQG transition amplitude
The covariant formulation of LQG defines transition amplitudes between quantum states of the geometry, in suitable truncations of the numbers of degrees of freedom. The number of degrees of freedom can be truncated by approximating the 3d geometry of the boundary by means of a cellular decomposition. The two skeleton of the dual of the decomposition form a graph and the truncated variables can be taken to be $SU(2)$ group elements $U_{l}$ on the links $l$ of this graph. The quantum states of the geometry can then be taken to be functions of the $U_{l}$’s in the space $L_{2}[SU(2)^{L}/SU(2)^{N}]$, where $L$ and $N$ are respectively the numbers of links and nodes of $\Gamma$, as in lattice Yang Mills theory. The algebra formed by the group elements $U_{l}$ themselves and the Left invariant operators is the observable algebra, which has a natural interpretation in terms of functions of the discretized geometry. A large literature has developed a theory of coherent states for this algebra, representing semiclassical geometrical states.
@@
Considering instead the entire quantum region simplifies the topology. The boundary of the quantum region can be take to be as depicted in blue Figure 10. The tip, at the largest radius, is a two sphere, chosen outside the Planckian curvature region. The lowest (past) part of the boundary and the upper (future) one are both 3d balls. Each of this can be triangulated with 4 tetrahedra and the interior can be triangulated with two four-simplices sharing an internal tetrahedron, as depicted in Figure 11.
Figure 11: The graph of the spin network and the two complex of the spinfoam for the calculation of the tranisition.
@@
The numerical calculations have been performed using different techniques. Frisoni2023 has analised the simpler geometry, utilizing the sl2cfoam-next code for computing spinfoam transition amplitudes. A good introduction to this, and full references, is in Dona2022a . The code permits to explore the low-spin (small-mass) regime of the amplitude. On the other hand, Han2024 utilizes the complex saddle point method for computing spinfoam amplitudes, which permits to explore the high-spin (large-mass) regime Han2020 ; Han2020 ; Han2023 ; Han2024a . A common result of all these investigations (that contradicts some previous expectations) is that the transition amplitude gives a transition probability proportional to
@@
Notice that this is indeed non-analytic in the Planck constant. Therefore it is a genuine non perturbative result that cannot be obtained in conventional perturbation theory expanding around a classical solution. The exponential is characteristic of quantum tunnelling phenomena. The transition, in fact, is forbidden classically (classically the geometry evolves into a singularity) and is characteristic quantum tunnelling effect.
@@
Anticipating the discussion about dissipative effects, we can see that toward the end of the Hawking radiation, $m$ decreases and approaches $m_{P}$. At this stage the transition becomes increasingly probable, going to probability unit when the mass becomes actually Planckian.

### v4

@@ I Introduction
Three aspects of this scenario are particularly appealing. (a) It offers a natural solution to the black hole information ‘paradox’. (b) It is in principle, and perhaps even in practice, directly testable Perez2023 . And (c) it provides a natural candidate for dark matter that does not require any new physics (such as new fields, particles, or modifications of the field equations): just general relativity and its quantum properties Rovelli2018f . This scenario is compatible with and possibly corroborating the general idea that primordial black holes could play a key role for explaining dark matter, see carr_primordial_2020 ; green_primordial_2021 .
@@
In the first part, we discuss the bounce of a collapsing matter distribution (Section II.1), namely the ‘Planck Star’ phenomenon, and the external metric of the global black hole dynamics (Section II.2). A side section discusses the counter-intuitive classical geometry of white holes, because this play a role in global picture (Section II.6). This part ends presenting the current state of the calculations of the black-to-white transition probability Christodoulou2018 .
@@ I.1 A sketch of the scenario
Perhaps the most interesting discovery of the last decades is the abundance and variety of the astrophysical objects we now call black holes. Current direct and indirect observations of these objects are all well accounted for by classical GR. But classical GR is insufficient to account for the phenomena that happen in their high-curvature regions, where genuine quantum gravity effects are most likely non-negligible. Therefore GR does not provide a reliable global picture of the dynamics of these objects.
@@
The internal geometry of these objects is a remarkably precise realization, fully consistent with classical GR, of the ‘cornucopia’ intuition considered in the 1980’s and later abandoned on the basis of arguments that (as we shall discuss below) are no longer cogent Banks1992b ; Banks1993 .
There is a well-known instability affecting eternal and macroscopic white holes: they can easily fall into a black hole as in the process described by the classical Kruskal geometry. On the other hand, there are arguments indicating that a Plank-mass white hole can be stabilized by quantum gravity Rovelli2018f , by shifting it into a quantum superposition of black and white geometries. As we show below, the key of this stabilization is the LQG area gap, because of which there is likely no black (or white) hole with a mass smaller than the Planck mass.
Hence the scenario predicts the existence of quasi-stable remnants of a known mass, determined by the LQG area gap. A large number of these, of primordial or possibly even pre-big-bang origin, would behave precisely as dark matter. The possibility of direct detection of these objects has been explored in Perez2023 and we review it below.
Once quantum gravity phenomena are taken into account, the notion itself of horizon is necessarily different than the one relevant for the classical theory. Event horizons lose their relevance, because they are likely not to form at all. The Hawking radiation drags the dynamical evolution of the horizon into regions where the Einstein equations do not hold anymore. Accordingly, the discussion on the so called black-hole information paradox needs to be reconsidered.
@@
While the scenario illustrated takes fully into account quantum gravitational phenomena in the high curvature region, it assumes that classical GR and quantum field theory on curved spacetimes are good approximations outside these regions. Contrary to what sometimes claimed, GR and quantum field theory taken together, yield no inconsistencies in the entire region where quantum gravity can be neglected. As argued in detail below, the common claim that inconsistencies appear already at the Page time depends on strong versions of the holographic hypothesis Rovelli2017a , or from confusion between ADM mass and Bondi mass, between thermodynamic and von Neumann entropies, or between event and dynamical horizons. On the contrary, the scenario considered here is only based on conventional GR and conventional quantum ideas.
@@ I.2 The domain of validity of classical gravity
The region of the onset of quantum gravitational phenomena can thus be organized into three sub-regions Dambrosio2021 (see Figure 1 ):
@@
which is colossal for a macroscopic black hole. Notice that the locus of the onset of quantum gravity is not ‘at a point’, but rather ‘at a time’. It is not ‘somewhere’ but rather, ‘at some time’, which, in relativistic terms, means in a set of space-like related regions.
The fact that the various onsets of quantum gravity are spatially related implies by locality that what happens in a region of this locus is causally disconnected, hence independent from what happens in other regions. To imagine that quantum gravity could violate this macroscopic causal disconnection is in gross contradiction with what we know about reality, not supported or suggest by anything we know, as far as we can see. To understand the physics of the end of a black hole, we have to understand the quantum evolution of each of these three regions independently.
@@ II.1 Planck stars
Let us start by studying what happens in the interior of a collapsing star, after it enters its horizon, when it reaches Planckian density. The simplest modelling of a collapsing matter distribution is provided by the Oppenheimer-Snyder model Oppenheimer:1939ue : a spherically symmetric pressure-less homogeneous matter distribution free-falling under its own weight. Assuming the matter distribution to start at rest in the distant past, the metric inside a such a matter distribution distribution can be written in co-moving and proper time coordinates $(T,R)$ as
@@
where the over-dot means differentiation with respect to $T$, which coincides with the proper time. Eq. (6) can be solved:
@@
where $A=3/(2\pi\rho_{c})$ is a parameter of Planckian value. In units where $G=c=1$, the constant is $A\sim\hbar\sim m_{Pl}^{2}$. Notice that $a(T)$ is positive for the whole range $T\in[-\infty,\infty]$: it decreases for $T<0$, reaches a minimum $a_{0}=\sqrt[3]{Am/2}$ for $T=0$ and then increases for $T>0$. See Fig. 2. This is the characteristic bounce of loop quantum cosmology. This result shows that the line element in eq. (5) is well defined everywhere, without singular collapse: the star reaches a maximal density, where it is called a ‘Planck star’, and then bounces and expands.
@@
To get a feeling for the physics of the bounce, we can rewrite (8) in the form
@@ II.2 Black-to-white transition
The maximal extension of this geometry has (i) two asymptotic regions and (ii) a time-like singularity in a high curvature region. These two features make it an unlikely candidate for describing the actual physics of realistic black holes. But a surprising result on spherically symmetric solutions of the Einstein’s field equations, solves both these difficulties. This is illustrated in the next section.
@@ II.3 The exterior metric
The Birkhoff’s theorem is sometimes presented as stating that the only spherically symmetric spacetime compatible with the vacuum Einstein’s field equations is the Kruskal geometry. At a sufficient distance from a black hole, we expect quantum effects to be negligible and therefore the theorem to hold. Hence we might expect that whatever happens where quantum effects are negligible must form a subset of a Kruskal geometry. In the Kruskal metric, there is an anti-trapped region, but it is before, not after the trapped region. Hence there seems no way to join the metric described above to a single external asymptotic region. But the above is wrong. The reason is that Birkhoff’s theorem is local, not global: it states that a spherically symmetric solution of the vacuum Einstein field equations is locally, and not necessarily globally, isometric to the Kruskal metric. This seems a red herring, but has momentous consequences for understanding the dynamics of quantum black holes.
@@
Figure 4: The spacetime discovered in Haggard2014 , describing the collapse of a null shell into a black hole and its bounce out as a white hole (left); and the way it can be locally isometrically mapped onto the Kruskal spacetime (right). The pink region is an exact solution of the classical Einstein equations.
@@
This same idea solves the two difficulties mentioned at the end of last section, providing a spacetime with a single asymptotic region and no singularities for the bouncing star. This can be constructed by cutting and pasting the maximal extension of the metric defined by eqs. (12) and (15), as follows.
@@
Next, delete the entire spacetime region enclosed within the blue line, and paste its two spacelike portions as in Figure 6. This is clearly possible since they are both constant Schwarzschild time surfaces.
@@
The resulting spacetime has a single asymptotic region, is everywhere locally isomorphic to the maximal extension of the metric (12-15), and has a hole, depicted as the $B$ diamond of Figure 6. In Han:2023wxg it is shown that the $B$ region admits a regular Lorentzian metric compatible with rest of the spacetime geometry. The $B$ region and its physics will be discussed later on. For the moment, let us focus on the rest of the spacetime.
@@
The first term on the right-hand side can be interpreted as the back and forth travel time for the light from the observer to the hole. The second term is the standard relativistic correction. The last term is independent of the radius and represents a property of the transition itself. The quantity $\delta$ can be measured by distant observers by measuring $T$ at different values of $R$. A little geometry shows that it depends on the choice of $\beta_{L}$ as follows: $r=2m+\delta$ is the radius at which the constant Schwarzschild-time surface passing by $\beta_{L}$ intersects the surface of the falling star. This is clearly very close to the collapse (the point where the surface enters entirely in its horizon), if the lifetime of the hole is long. Thus
@@
can be taken as a definition if the duration of the quantum process, as observed from infinity. In our construction, it is determined by how the geometry has been constructed by cutting and pasting. Operationally, it is the observed duration of the process from very large distance. Physically, it is determined by the quantum theory of the geometry of the $B$ region, to which we now turn.
@@ II.4 The Boundary Region
To address this question we need to step back and think more precisely what we are doing. First, let us remember once more that for the moment we are disregarding the dissipative effects that break time reversal invariance. These will be studied in the next section. Then, consider the fact that we are describing a quantum process. There are different ways of describing quantum processes. In some regimes, these can be described as corrections to the classical equations of motion. This is not viable for the quantum transition of the horizon, which is a non-perturbative process. One cannot describe quantum tunnelling as a simple $\hbar$ correctios to classical trajectories. Tunnelling is non-analytical in $\hbar$.
Another possibility is to describe the evolution of a quantum state. Quantum states that approximate classical configurations, such as coherent states, spread. We could therefore consider the full quantum state of the geometry after the actual quantum process as a quantum state, namely a quantum superposition of geometries. This is analogous to describing nuclear radioactivity in terms of a wave function of the emitted particle widely spread in space and time. Correct, but incomplete. What we observe in nuclear radioactivity is not the wave function of the emitted particle, widely spread in space and time. Rather, it is a specific space and time location where the particle is detected, say by a Geiger counter.
@@
That is, we can compute the transition amplitude for the geometry to make a quantum transition from its value in the past boundary of $B$ to its future boundary (as opposed to not yet making it, as in radioactivity). Notice that the geometry (both intrinsic and extrinsic) of these two boundaries is known from the construction of the above section. We have simply to compute the LQG transition amplitude between quantum states approximating these geometries.
The covariant formulation of LQG is precisely formulated to do this calculation, using approximations provided by truncation of the degrees of freedom. The calculation has been performed, under some drastic simplifications by Christodoulou and D’Ambrosio using analytical techniques in Christodoulou2018 ; Dambrosio2021 ; Christodoulou2023 . The result has been confirmed numerically in Frisoni2023 and Han2024 . Numerical calculations are been currently developed to go beyond these approximations. The way these calculations are performed is summarized in the next section.
@@ II.5 The LQG transition amplitude
The covariant formulation of LQG defines transition amplitudes between quantum states of the geometry, in suitable truncations of the numbers of degrees of freedom. The number of degrees of freedom can be truncated by approximating the 3d geometry of the boundary by means of a cellular decomposition. The two-skeleton of the dual of the decomposition form a graph and the truncated variables can be taken to be $SU(2)$ group elements $U_{l}$ on the links $l$ of this graph. The quantum states of the geometry can then be taken to be functions of the $U_{l}$’s in the space $L_{2}[SU(2)^{L}/SU(2)^{N}]$, where $L$ and $N$ are respectively the numbers of links and nodes of $\Gamma$, as in lattice Yang Mills theory. The algebra formed by the group elements $U_{l}$ themselves and the left-invariant operators is the observable algebra, which has a natural interpretation in terms of functions of the discretized geometry. A large literature has developed a theory of coherent states for this algebra, representing semi-classical geometrical states.
@@
Considering instead the entire quantum region simplifies the topology. The boundary of the quantum region can be take to be as depicted in blue Figure 10. The tip, at the largest radius, is a two-sphere, chosen outside the Planckian curvature region. The lowest (past) part of the boundary and the upper (future) one are both 3d balls. Each of these can be triangulated with four tetrahedra and the interior can be triangulated with two four-simplices sharing an internal tetrahedron, as depicted in Figure 11.
Figure 11: The graph of the spin network and the two-complex of the spinfoam for the calculation of the transition.
@@
The numerical calculations have been performed using different techniques. Frisoni2023 has analyzed the simpler geometry, utilizing the sl2cfoam-next code for computing spinfoam transition amplitudes. A good introduction to this, and full references, is in Dona2022a . The code can explore the low-spin (small-mass) regime of the amplitude. On the other hand, Han2024 utilizes the complex saddle point method for computing spinfoam amplitudes, which can explore the high-spin (large-mass) regime Han2020 ; Han2020 ; Han2023 ; Han2024a . A common result of all these investigations (that contradicts some previous expectations) is that the transition amplitude gives a transition probability proportional to
@@
Notice that this is indeed non-analytic in the Planck constant. Therefore it is a genuine non-perturbative result that cannot be obtained in conventional perturbation theory expanding around a classical solution. The exponential is characteristic of quantum tunnelling phenomena. The transition, in fact, is forbidden classically (classically the geometry evolves into a singularity) and is a characteristic quantum tunnelling effect.
@@
Anticipating the discussion about dissipative effects, we can see that toward the end of the Hawking radiation, $m$ decreases and approaches $m_{P}$. At this stage the transition becomes increasingly probable, going to probability unity when the mass becomes actually Planckian.